OpenStudy (anonymous):

A sport car moving at constant speed travels 110m in 5.0 s. if it then brakes and comes to a stop in 4.0 s, what is the magnitude of its acceleration in m/s2, and in g’s (g=9.80 m/s2)? 2

OpenStudy (anonymous):

It was initially moving with velocity,u=110/5=22m/s. Final velocity=0 Time taken to retard= 4s v=u+at gives a=-5.5m/s In terms of g, a=5.5/9.8 times g.

OpenStudy (anonymous):

that is what I came up with thanks